OpenStudy (aravindg):

A cylindrical tank of height 0.4 m is open at the top and has a diameter 0.16 m.Water is filled in it upto a height of 0.16 m .Calculate the time taken to empty the tank through a hole of radius 5x10^-3m in its bottom?

OpenStudy (aravindg):

@Vincent-Lyon.Fr , @ash2326 , @waterineyes

OpenStudy (vincent-lyon.fr):

First, use Torricelli's law to determine speed of water in the hole and rate of flow. \(v=\sqrt{2gh}\)

OpenStudy (aravindg):

then?

OpenStudy (vincent-lyon.fr):

Is the answer 45.80 s ?

OpenStudy (aravindg):

i didnt get tht :O

OpenStudy (aravindg):

i got 2.8 :P

OpenStudy (vincent-lyon.fr):

How?

OpenStudy (aravindg):

wait

OpenStudy (anonymous):

h is variable

OpenStudy (aravindg):

hmm

OpenStudy (aravindg):

then how ?

OpenStudy (vincent-lyon.fr):

Total volume is 3.217 L, right?

OpenStudy (aravindg):

i got 3.14x10^-4

OpenStudy (vincent-lyon.fr):

Try again. radius is 0.08m and height is 0.16m

OpenStudy (aravindg):

3.21x10^-3

OpenStudy (vincent-lyon.fr):

ok

OpenStudy (aravindg):

so?

OpenStudy (vincent-lyon.fr):

now escape velocity is \(v=\sqrt{2gh}\)

OpenStudy (vincent-lyon.fr):

and discharge is v multiplied by area of hole.

OpenStudy (aravindg):

v is velocity ryt?

OpenStudy (vincent-lyon.fr):

yep

OpenStudy (aravindg):

so how do we bring in volume?

OpenStudy (vincent-lyon.fr):

time taken is: volume / discharge you should find 22.9s

OpenStudy (aravindg):

hmm

OpenStudy (aravindg):

but isnt h varying?

OpenStudy (vincent-lyon.fr):

Yes, but we can get round this difficulty

OpenStudy (vincent-lyon.fr):

Do you find the same values?

OpenStudy (aravindg):

how/?

OpenStudy (vincent-lyon.fr):

initial escape velocity is v=root(2g hmax) discharge is v multiplied by area of hole. time taken is: volume / discharge you should find 22.9s This is the time it would take to empty the tank if initial discharge was constant. Real duration, when h varies, is twice that time.